### v1

# The Serial Scaling Hypothesis

Abstract
While machine learning has advanced through massive parallelization, we identify a critical blind spot: some problems are fundamentally sequential. These "inherently serial" problems—from mathematical reasoning to physical simulations to sequential decision-making—require dependent computational steps that cannot be parallelized. Drawing from complexity theory, we formalize this distinction and demonstrate that current parallel-centric architectures face fundamental limitations on such tasks. We argue that recognizing the serial nature of computation holds profound implications on machine learning, model design, hardware development. As AI tackles increasingly complex reasoning, deliberately scaling serial computation—not just parallel computation—is essential for continued progress.
1 Introduction
Scaling up machine learning has driven remarkable progress [1, 44, 26, 49, 55], but not all scaling is created equal. Although the field has focused on increasing parallel computation [23, 17, 87, 102, 46]—with bigger GPUs, wider models, and more data—some problems stubbornly resist these advances [2, 51, 59, 71, 69, 16]. For a class of tasks, only increasing serial computation—allowing models to perform more sequential steps—yields further progress.
The recognition that some problems fundamentally require deep or sequential models has remained mostly a theoretical curiosity [115, 72, 13]. Only recently has there been a growing recognition of scaling test-time computation, in contrast to train-time computation [105, 82, 128]. Yet this dichotomy still overlooks the role of parallel vs. serial computation which is central to this paper. In practice, both parameter counts and computation FLOPs are largely treated as single numbers, with no distinction between width (parallel) and depth (serial computation) [49, 44, 82, 105].
Consider Sudoku—a number-placement puzzle requiring each number 1–9 to appear once per row, column, and subgrid—as a parable (Figure 1). Easy puzzles can be solved by filling in many blanks independently, in parallel. Hard puzzles, however, require a long chain of dependent reasoning: each blank depends on the others. No algorithm can shortcut the process. This distinction—between problems that are “wide” (parallel) and those that are “deep” (inherently serial)—is fundamental; it dictates how we design efficient models, yet it is underappreciated in machine learning.
We propose the Serial Scaling Hypothesis: For many important ML problems, especially those involving complex reasoning, planning, or the evolution of interacting systems, increasing parallel computation alone is insufficient. Progress requires scaling the amount of serial computation.
This hypothesis is motivated by both theory and practice. Complexity theory shows that some problems cannot be efficiently parallelized. Empirically, tasks such as hard Sudoku, mathematical reasoning, and sequential decision-making benefit disproportionately from models that can perform more serial steps (e.g., deeper networks, longer chains of thought, more search iterations).
Why does this matter for machine learning? As we push toward more challenging tasks—advanced reasoning, physical simulations, planning, and scientific discovery—we encounter problems that parallel architectures (like Transformers and diffusion models) cannot efficiently solve. Recognizing the limits of parallel scaling and the necessity of serial computation has several implications:
  • Model design: Should we revisit architectures that allow for deeper or more sequential computation?
  • Hardware: Is it time to invest in faster, lower-latency processors, not just more parallel ones?
  • Evaluation: Should we measure and report serial compute separately from total compute?
Contributions. We (i) formulate the Serial Scaling Hypothesis and connect it to complexity theory, (ii) survey a range of problems—cellular automata, many-body physics, sequential decision-making, and advanced reasoning—where serial computation is provably or empirically essential, (iii) analyze the limitations of modern parallel-centric machine learning models on these tasks, and (iv) discuss implications for future model and hardware design, with practical recommendations.
1.1 A Parable: Easy vs. Hard Sudoku
Figure 1: The four circled blanks in the easy Sudoku have no interdependence and can be filled in parallel. The blanks in a hard Sudoku puzzle are densely interdependent, such that one blank cannot be filled conclusively until one has worked out a long chain of consequences for many others.
Not all hard problems are hard in the same way. An easy Sudoku puzzle can be solved by filling in many blanks independently—each can be completed in parallel. In contrast, a hard Sudoku requires a long chain of dependent reasoning: each blank depends on the others, and the solution must be built step by step. No algorithm exists that can shortcut this process; only more sequential computation—more "depth"—helps. While a large number of easy puzzles may take the same amount of computation as a single hard puzzle, yet one can be significantly accelerated by more processors, while the other cannot. This distinction between “wide” (parallelizable) and “deep” (inherently serial) computation is central to our thesis. Many tasks in machine learning, from advanced reasoning to planning, share this serial character: they require a sequence of dependent steps, not just more parallel computation.
1.2 A Brief History of Scaling
Serial Parallel
Scaling Depth Width
Architecture RNN Transformer
Algorithm RL Imitation learn.
Hardware CPU GPU
Table 1: Examples of serial and parallel approaches.
The success of modern machine learning has been driven by scaling—bigger models, more data, and especially more parallel compute. Hardware has shifted from CPUs to massively parallel GPUs; architectures have moved from RNNs to highly parallelizable Transformers; and algorithms increasingly exploit parallelism for efficiency.
However, this focus on parallel scaling has a blind spot: it assumes that parallel and serial computation are interchangeable. In reality, for many problems, only increasing serial compute—allowing models to perform more sequential steps—yields further progress. This is especially true for tasks with inherent temporal or causal dependencies. As we show in this paper, recognizing the limits of parallel scaling and the necessity of serial computation is crucial for the next phase of progress in machine learning.
2 Machine Learning from the Serial Perspective
Figure 2: . (A) A decision problem has a variable-size input and a fixed-size output (e.g., “yes”/“no”). (B) A serial problem requires deeper or more steps as the problem size grows. Examples of serial problems are: (C) Cellular automaton: takes the initial state as input and outputs a discrete value of the row $N$ at cell $i$ for $i\in\{1,\dots,2N-1\}$. (D) Many-body mechanics: takes initial positions and momenta of each particle with time $T$ as inputs and outputs the particle locations at time $T$ in a limited-precision space. (E) Math QA: takes a question as input and outputs the answer autoregressively, with each output from a fixed set of possibilities.
We first formalize inherently serial problems (Section 2.1), drawing from complexity theory to show they are not only theoretically valid but also pervasive in machine learning domains (Section 2.2). We then examine why predominant ML models struggle with these tasks (Section 2.3). Finally, we discuss broader implications for both software and hardware design (Section 2.4).
2.1 Complexity Framework for Inherently Serial Problems
We focus our attention on decision problems which take in $N$ input tokens and output either “yes” or “no” as depicted in Figure 2(A). While this might seem limited, it is incredibly general. Any problem with discrete output can be framed as a decision problem by asking multiple yes/no questions to cumulatively reconstruct the final output. Figure 2 (C–E) show how cellular automata, many-body mechanics, and open-ended question answering can be represented in this way.
Figure 3: The complexity classes are nested as $\mathsf{TC}^{0}\subseteq\mathsf{TC}^{1}\subseteq\dots\subseteq\mathsf{TC}\subseteq\mathsf{P}$. Each containment is widely believed to be strict. Problems in $\mathsf{TC}$ are considered parallel, while problems outside are inherently serial.
As our focus is on recent machine learning developments, i.e., deep neural networks, we adopt the complexity class $\mathsf{TC}$ (Threshold Circuits) which has strong connections to neural networks to formally distinguish between serial and parallel problems.
Definition 2.1 (Informal, see Appendix A).
A problem belongs to the class ($\sf{L}$-uniform) $\mathsf{TC}^{i}$ if it can be solved by a family of Boolean circuits—one for each problem size $N$—with $\mathsf{poly}(N)$ width and $\mathcal{O}(\log^{i}N)$ depth, using gates such as AND, OR, NOT, and MAJORITY. Here, $\sf{L}$-uniformity means that there exists a single program that would output the schematics of the $N$-th circuit using only $\mathcal{O}(\log N)$ space during the computation. The class $\mathsf{TC}$ is defined as the union of all such $\mathsf{TC}^{i}$ for $i\geq 0$.
Intuitively, a problem belongs to $\mathsf{TC}$ if and only if it can be solved by an MLP with polynomial width and polylog depth, where polylog denotes $\sf{poly}(\log N)$ [85]. A problem belongs to $\mathsf{TC}^{0}$ if and only if it can be solved by a constant-depth MLP. We consider $\mathsf{TC}$ problems as parallel problems since there are sublinear-depth MLPs that solve them.
While that seems contradictory with the universal approximation theorem [22, 45] which argues that a 3-layer MLP with an unbounded width may approximate any continuous function, it does not. When the width is bounded as polynomial, the problems that the MLP can solve are much more restricted to only in $\mathsf{TC}^{0}$, a class of problems that fixed-depth MLPs can solve.
It is a common phenomenon in computational complexity that one can make a neural network shallower at the cost of exponentially larger width [118, 40, 101, 124, 81, 28, 20, 115, 72]. However, such exponential-sized networks are intractable in both memory and computation. We consider them as inefficient solutions in this paper. This exponential depth-width trade-off underlines the importance of characterizing neural networks in terms of depth, which motivates the Serial Scaling Hypothesis.
To formally characterize the problems in terms of depth, we have the following definition:
Definition 2.2.
A problem $\mathcal{P}$ is parallel if $\mathcal{P}\in\mathsf{TC}$; otherwise, it is inherently serial.
Assumption 2.3.
We adopt the widely held belief that $\mathsf{TC}\subsetneq\mathsf{P}$ [34, Ch. 5]—that is, some polynomial-time problems, e.g., $\mathsf{P}$-complete ones, are inherently serial [34, Ch. 8].
Importantly, labeling a problem inherently serial does not preclude the use of parallel computation. It is possible that a serial problem might have most of its problem instances that are parallelizable, and parallel solutions, such as sublinear-depth MLPs, may obtain a respectable average accuracy. However, solving such a problem exactly would still require a genuinely serial solution, such as linear-depth MLPs or more.
2.2 Real-world problems are likely inherently serial
The existence of inherently serial problems does not automatically imply their practical relevance. A natural question arises: “Are these problems likely to appear in real-world applications?” The answer is: Yes—particularly if one works in domains involving the need for physical simulation or reinforcement learning. But even outside these areas, it remains highly likely that one will encounter inherently serial problems. Surprisingly, even tasks that appear simple and for which we have a polynomial-time computer program can exhibit inherent seriality that resists parallelization. [^1] Certain problems are believed to be inherently serial despite being solvable in polynomial time, due to the strong belief that $\mathsf{TC}\neq\mathsf{P}$ [34, Ch. 5].
Drawing from both empirical observations and theoretical arguments presented in the following Section 4, we propose the following complexity-theoretic hypothesis:
Hypothesis 1 The following problems—cellular automata evolution, many-body mechanics, sequential decision problems, mathematical question answering—are inherently serial (i.e., not contained in the complexity class $\mathsf{TC}$).
Problems outside the class $\mathsf{TC}$ (see Definition 2.1) exhibit fundamental computational dependencies: the outcome of intermediate steps directly influences subsequent steps in ways that cannot be shortcut without compromising correctness. This intrinsic structure precludes flattening the execution into a parallel form without loss of fidelity.
2.3 Are our models capable of solving inherently serial problems?
Knowing how relevant inherently serial problems are, the crucial question is: “Are our models also serial?” The answer is they may not be.
The predominant computational architectures adopted in our community excel in parallel computation efficiency. Transformers and state-space models (SSM) [36, 35] are themselves “sequence” models, yet their computation is parallelizable. Unless used with autoregressive inference, which limits their throughput to one token at a time, a Transformer or SSM layer can ingest the whole $N$ input tokens concurrently and produce the output.
Based on theoretical evidence provided in Appendix B, this seeming efficiency comes at a cost:
Theorem 2.4 (Informal [73, 15, 14, 74]).
Any computation performed by a standard feedforward (non-recurrent) network, a fixed-precision Transformer, or a linear state-space model of polynomial size and constant depth can be simulated by a uniform threshold circuit of constant depth—i.e., $\mathsf{TC}^{0}$.
Membership in $\mathsf{TC}^{0}$ means a model’s computation can be executed in a constant number of parallel steps, with each layer flattened into a threshold circuit, allowing all input tokens to be processed simultaneously—not relying on long chains of sequential computation.
In contrast, truly serial models—such as recurrent neural networks (RNNs), repeating layers, and autoregressive Chain-of-Thought (CoT)—fall outside $\mathsf{TC}$, reflecting their ability to capture dependencies that unfold over the input tokens.
It is important to note that a parallel architecture doesn’t necessarily mean a parallel model. As a model is a combination of architecture and inference method, a parallel architecture such as a Transformer, when used with autoregressive Chain-of-Thought inference, becomes a serial model. Section 3.1 provides the discussion in more detail.
Method Parallel? Solve serial problem?
MLPs
Transformers
SSMs and Mamba
RNNs
Repeating layers
Chain-of-Thought
Diffusion models ($\sf{TC}^{0}$ backbone) 22footnotemark: 2
Table 2: Parallelizable models are limited to parallel problems. Only non-parallelizable models may solve inherently serial problems.
footnotetext: While diffusion models are not yet parallelizable, given their $\mathsf{TC}^{0}$ capabilities, there might exist a parallelization algorithm for diffusion models.
Diffusion models cannot solve inherently serial problems. The surprising finding that is first proven here is that diffusion models—which appear non-parallelizable due to their iterative nature—are incapable of solving inherently serial problems. We demonstrate theoretically in Section 5 that diffusion models [42, 107, 106] with a fixed-depth $\mathsf{TC}^{0}$ backbone network still lie in $\mathsf{TC}^{0}$, i.e., cannot solve inherently serial problems, even with infinite diffusion steps. We provide a summary of models’ theoretical capabilities in Table 2.
Under the Section 2.2, we conclude a result that we believe is particularly relevant and accessible to practitioners in the machine learning community:
Key Limitation of Modern Machine Learning Models
Theorem 2.5 (Informal).
Transformer-based, structured state-space models (SSMs), and diffusion models with $\mathsf{TC}^{0}$ backbones are provably incapable of solving general instances of problems such as cellular automata evolution, many-body mechanics, sequential decision-making, and mathematical question answering.
By connecting the model’s computation characteristics with the inherently serial nature of the key problems, this theorem highlights the fundamental insufficiency of the predominant non-serial models we use and the inherently serial problems we aim to solve.
Limitations. (1) Some instances of inherently serial problems may be parallelizable or even trivial; only in the general case do they exhibit seriality. (2) For exceedingly hard problems in $\mathsf{NP}\setminus\mathsf{P}$ and beyond, serial computation may not be the dominant bottleneck due to the overwhelming exponential cost. Therefore, our position is most applicable to problems of practical real-world difficulty.
2.4 Implications of The Serial Scaling Hypothesis
For machine learning practitioners. Many important real-world problems—such as cellular automata, many-body mechanics, and sequential decision-making—are inherently serial. Solving them requires sufficient serial computation unless a model is exponentially wide. But in practice, we can afford neither exponentially wide models nor exponentially large datasets. This mismatch—solving inherently serial problems using shallow or parallel models—may help explain why many ML models only work well on examples similar to their training data [117, 135, 89, 60, 68, 136].
We hypothesize that models during training learn many algorithms for each problem. Some take constant time to run, essentially “memorization” and “shallow pattern-matching.” Some take a short time to run and correspond to “fast heuristics.” Some take longer to run but produce correct, general solutions. However, at inference time, if the model lacks enough serial computation, it defaults to the fastest routines—those that fail to generalize beyond the training distribution.
Recent evidence suggests that pretrained LLMs are already capable of both short-form and long-form solutions. However, only after RL fine-tuning is a longer and often more general solution elicited [123, 134, 100]. We thus advise that for models to generalize, they should be given sufficient serial computation at inference time.
For model designers. To solve these challenging real-world problems efficiently, future models may need to incorporate recurrent structures to increase their serial computation, in addition to the current predominantly parallel designs. It is important to note that such designs may come with trainability concerns. For instance, adding recurrence or depth often increases the gradient variance [8, 86] and the L-Lipschitzness of the model [7, 30], which cumulatively make the model harder to train. This further requires improvements in training techniques.
For hardware designers. The need for more serial computation is fundamental and cannot be substituted by parallel computation. Future progress in machine learning depends on improving low-latency, sequential processing. This is a call to potentially revive interest in CPUs or specialized processors that tightly integrate memory and compute to reduce data movement overhead [48, 50]—while still leveraging parallelism within each step, as in GPUs.
3 Potential Misconceptions
3.1 Architecture vs. Inference Methods
One may be confused by the statement that “Transformers are not serial,” followed by “Chain of Thought is serial.” This appears to be a contradiction. It is not, though it may appear so due to our omission of words. The first statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, but only run it for $\mathcal{O}(1)$ forward passes, then you cannot solve inherently serial problem-families.” Similarly, the second statement really should be “If you fix a Transformer’s parameters and give it problems of size $\mathcal{O}(n)$, and also run it for $\mathcal{O}(n)$ forward passes, then you can solve inherently serial problem-families.”
This is a special case of a general phenomenon: people confuse architecture and inference methods, meshing them all together like two lumps of clay gently melding under the desert sun. This is unfortunate but understandable because, in practice, architecture and inference couple together strongly, and this confusion is also encouraged by things like “architecture–deployment co-design.” The fact is that the abstractions leak. How you use a model for inference depends on its architecture. How you design a new model architecture depends on what you expect its inference method to be. This is a good thing, but it may cause people to confuse architecture and inference.
Fundamentally, we consider an “architecture” to be a class, and its instances are “atoms of computation,” and this is the intuition that the reader should keep in mind. An atom has mass and volume. An atom can be copied and put into many molecules. What an atom cannot do is be divided or changed. Similarly, an instance of the “Transformer architecture” is an “atomic” Transformer. This atomic Transformer does have subatomic structures, such as its individual layers. However, since we are not performing complicated model surgeries, such as taking off its head, reading out the tensor from its middle layers, etc., we can assume it is atomic. The least you can do with this atomic Transformer is perform one single forward pass, and this has a definite computational cost. It may be 48 layers or 96 layers, but the cost is definite.
Still, even though the atomic Transformer is unchanging, it can be used in many molecules, and the molecules can change. It can be used in a Chain of Thought molecule, and this molecule would grow like a polymer, taking on a length of $\mathcal{O}(n)$ in reaction to a length $\mathcal{O}(n)$ input.
In short, a single architecture may be used on its own in a fixed number of forward passes, or a number of forward passes growing linearly with the problem size, or in a tree search, a beam search, etc. Inference methods are separate from model architecture. In this paper, we discuss both architectures and inference methods. We have found that Transformers, SSMs, and diffusion models are all not serial when the inference method is just a single forward pass, but can become serial when the inference method allows for more serial compute, such as Chain of Thought inference.
3.2 SSM is not RNN
We class SSMs as not inherently serial, while RNNs as inherently serial. This is not a mistake, and in fact, a very valuable illustration of how an architecture can appear serial but is not.
First, in what sense is RNN serial, while a Transformer is not serial? Recall the title of the original paper on Transformers: “Attention is all you need” [119]. What did they mean by “all you need”? What was removed? Back in 2017, the standard sequence transduction model was a pair of RNNs connected in the middle by the attention mechanism. The encoder RNN processes the input sequence one token at a time, then the decoder RNN produces the output sequence, again one token at a time. The key problem is that, while both an RNN and a Transformer must decode tokens one-by-one, an RNN must encode tokens also one-by-one, whereas a Transformer can encode tokens all-in-one-go. The animating spirit of the Transformer architecture was to remove everything that stands in the way of all-in-one-go, but preserve everything that allows RNNs to work so well.
So they removed recurrence and preserved attention. Their title meant “attention is all you need (and recurrence is not).” This is exactly what our paper is arguing against: attention is not all you need, and inherently serial processing is necessary. This may be performed by recurrence or some other architecture, but it cannot be avoided.
So now we come to SSM. SSM is not RNN because it is not recurrent in an inherently serial way. It was designed with the same spirit as the Transformer and suffers the same issue.
Concretely, consider the problem of ingesting an input sequence: $x_{1},x_{2},\dots,x_{n}$, and the task is to output the next token $x_{n+1}$. For a Transformer with 96 layers to accomplish the task, it does so by 96 sequential steps, each step is very wide but not deep. In contrast, an RNN with 96 layers must process $x_{1}$ first, then $x_{2}$, and so on. It requires $96n$ sequential steps that cannot be done in parallel. Suppose one attempts to do them “in parallel,” then there is an immediate problem: To run the RNN on $x_{2}$, one needs the internal states of the RNN just after it has processed $x_{3}$, and to do that, it needs to have processed $x_{2}$, etc. Abstractly, the operation of an RNN looks like:
$f_{\theta}(x_{n},f_{\theta}(x_{n-1},\dots,f_{\theta}(x_{2},f_{\theta}(x_{1},s_{0})))),$
where $s_{0}$ is the initial internal state.
In short, because the internal states of an RNN change nonlinearly and unpredictably, one cannot skip the steps.
The animating idea of SSM is precisely to remove the inherent seriality of RNN. The idea is that if the internal states change linearly and predictably, then one can skip the steps. Concretely, the operation of an SSM is of the form:
$\text{output}_{k}=f_{\theta}(M^{k-1}x_{1}+M^{k-2}x_{2}+\dots+M^{1}x_{k-1},x_{k})$
since its internal states change linearly in a data-independent way. This means that to run an SSM on input $x_{4}$, one does not need to know the internal state of the SSM after processing $x_{3}$. This allows it to be just as parallel as a Transformer, and thus just as lacking in seriality.
3.3 Training vs. Inference
Throughout most of the paper, the main contrast is the parallel vs. serial computation contrast. We do not mean this to be a contrast between the parallel phase of training vs. the autoregressive phase of inference, which is how modern GPT-like transformers are produced and used.
Concretely, consider initializing a language model with 96 layers. The training algorithm samples a chunk of text (32,768 tokens long) from the training corpus. The model then performs a single forward inference simultaneously on all tokens, such that it only takes 96 steps of time—one step per layer. Now, during inference, the model must then generate one token at a time, so the same 32,768 tokens would take 3,145,728 steps of time, which is a vast expansion.
Yet, this is not the main focus of the paper. The paper mostly talks about the theoretical and empirical results concerning learned and frozen models. The paper does talk about training, but only occasionally and informally, by sketching out intuitive analogies with inference, without rigorous theoretical justification.
This is not due to an intention to mislead. Indeed, we recognize that there is a parallel training vs. autoregressive inference contrast, and we have attempted to get some theoretical handle on this contrast. We failed. Because learning theory is extremely difficult, we could neither find theorems proven by people that came before us, nor prove theorems ourselves. Therefore, we stayed within inference. Fortunately, the parallel vs. serial contrast is already sharp even within inference, with both clear theoretical and empirical results, allowing us to write the paper.
We believe strongly that in the future, the parallel vs. serial contrast will also be shown for training, and leave this as work for the near-posterity.
4 Serial Problems
In this section, we introduce inherently serial problems. We begin by showing that even a simple evolving system (Section 4.1) exhibits this property. We then demonstrate that a system governed by physical laws (Section 4.2) also resists parallelization. We then show that inherently serial problems are quite generic, appearing throughout computational complexity theory (Section 4.3). Finally, we describe how sequential decision making (Section 4.4) and math QA (Section 4.5) are bottlenecked by serial computation.
4.1 Cellular Automata
Figure 4: Unpredictable patterns of Rule 110 cellular automaton and its rules. Given the top row, CA evolves to the bottom row, one by one, according to the rules.
We begin with the problem of predicting the outcome of cellular automata (CA) [95, 19], shown in Figure 2(C), as an example of a simple problem that is inherently serial. A CA consists of a regular grid of cells, each of which can be in one of finitely many states. The system evolves from an initial state in discrete time steps according to a local update rule that determines the new state of each cell based on the states of its neighboring cells. Despite their simplicity, CA can exhibit a wide spectrum of behaviors, ranging from completely predictable to extremely complex and seemingly random dynamics such as Rule 110 in Figure 4. The unpredictability is the first hint at the inherent seriality of this problem.
Cellular automata problem is inherently serial. The Rule 110 cellular automaton has been proven to be Turing-complete [127], meaning that there is an algorithm which converts a Turing machine to an initial state (top row) of the CA such that steps of the Turing machine are performed by iterations of the CA. Due to Turing-completeness, the problem of computing the state of a particular cell $x_{i}$ at row $N$ of Rule 110 CA is only solvable by simply working out the state evolution, one after another—without shortcuts [34, p. 58]. This is not unique for Rule 110. Many CA problems are known to be $\sf{P}$-complete [76], i.e., not efficiently parallelizable.
One might attempt to reduce serial computation by evolving across $k$ rows at once, as in the linear speedup theorem [84, Sec. 2.4]. However, this requires examining $12k+$ 1 adjacent cells, leading to $2^{2k+1}$ possible rules that must be stored and looked up—much larger than the original 8. While this reduces serial compute linearly, it increases parallel cost exponentially. This exponential depth-width tradeoff is a common phenomenon in computational complexity [118, 40, 101, 124, 81, 28, 20, 115, 72]. See  [47, Sec. 11.10] for a review.
What does this mean? A simple system evolving under 8 rules is sufficient to eliminate computational shortcuts, requiring serial solutions. Reducing the serial computation often comes at an exponential cost in parallel computation. Serial and parallel computations are not interchangeable.
4.2 Many-Body Mechanics
The previous section shows that inherent seriality arises from a simple evolution of a system governed by just 8 rules. Here, we examine a more realistic case: a many-body system under Newtonian mechanics. Consider $N$ particles moving in a multi-dimensional space under a force field or hard collisions. As shown in Figure 2(D), given the starting positions and momentum of particles at time $t=0$, predict the positions of the particles at some later time $t=T$ in a space with limited precision. [^3] Limited precision is an assumption used in practical scientific settings. With unlimited precision, the problem becomes $\mathsf{PSPACE}$-hard, i.e. intractable.
Many-body mechanics are inherently serial. The key argument is that Newtonian physics is rich enough to simulate any computation. Fredkin & Toffoli [32] recreate any Turing machine using nothing but billiard balls elastically reflecting off each other and walls—called a billiard-ball computer. Moore [75] proposes a similar recreation using smooth particle motion in a potential field in $\mathbb{R}^{3}$. Since the problem of simulating general Turing machines is inherently serial [34, p. 58], an algorithm that accurately simulates general physical systems can only be done step by step—except the special cases where the underlying Turing machine configurations are parallelizable by design.
Physical systems are often described as ODEs, which raises the question: “Can ordinary differential equations (ODEs) be parallelized over time?” The answer is: it depends. As a case study: Parareal algorithm [62] parallelizes the integration by applying a coarse solver to divide the ODE into multiple parts $[0,t_{1}],[t_{1},t_{2}],\dots,[t_{N-1},T]$—where each segment can be solved in parallel—followed by iterative corrections to reconcile the trajectories. Parareal may yield speedup if the coarse solver is sufficiently accurate—i.e., local gradient information reflects the global behavior—and may offer no wall-clock speed up otherwise [67, 33, Rmk. 4.7]. This leads to an intuitive rule of thumb: slow-changing ODEs tend to be parallelizable; fast-changing ODEs tend to be inherently serial.
Figure 5: To predict the next frame at time $T$, the input frames are given, however, the intermediate frames are effectively masked due to camera motion or occlusions.
Such seriality induced by object interactions is not exclusive to scientific simulations but also rooted in video prediction, where the task is to predict the next frame given a sequence of $N$ video frames. This setup is used to train large-scale models for content creation [133, 12, 130, 122, 114] and decision making [6, 29, 129].
Video prediction is inherently serial. This is because video frames are still rich enough to capture the mechanics of a 2D billiard ball computer. It takes two consecutive frames to infer the momentum of each ball, assuming unit mass. To predict the particles positions at time $T$, we mask the intermediate frames until time $T$, then the output is the frame at time $T$. Similarly in real-world videos, objects frequently move out of view due to camera motion or occlusions, effectively acting as masked frames, as illustrated in Figure 5. Thus, video prediction—in theory or practice—requires predicting object positions many steps into the future, which is inherently serial whenever a complex object interaction occurs with any regularity.
Inherent seriality calls for a model that tracks the evolving state of all objects—occluded ones included. If the model loses track of any object, it must serially compute its current position from the last known location. Not only taking long wall-clock time, it also makes the task unsuitable for parallel models such as Transformers.
What does this mean? Tasks involving the modeling of future dynamics of a system, including physical simulations and video prediction, likely require serial models to be efficiently solved. For video prediction, the video model may also need to maintain an evolving world state to keep the task tractable.
4.3 $\mathsf{P}$-complete Problems
The inherently serial problems given previously have the following commonality: Given a problem statement of size $\mathcal{O}(n)$, it is immediately clear how to construct a linear computational graph of length $\mathcal{O}(n)$, such that each node of the graph only depends on the previous one, and each step takes constant time. This is quite intuitive since this ought to be what being “serial” means.
In computational complexity theory, there is a large class of problems of this form, or reducible to a problem of this form: $\mathsf{P}$-complete problems. They have been studied in the theory of parallel computation. They serve to formalize the intuition that certain problems are efficient to solve on a single processor but impossible to solve much more efficiently with polynomially-many processors. The precise definition is in Appendix A.
Of the $\mathsf{P}$-complete problems, the most prototypical ones involve the evaluation of Boolean circuits, called the Circuit Value Problem (CVP) [56]. The CVP is quite robust, in that many variations on it are still $\mathsf{P}$-complete [34]. The CVP problem is: Given a Boolean circuit with a single output bit, specified as a list of logic gates and how they connect to each other, and an input binary string, what is the output bit?
The obvious solution is to perform a topological sort on the graph of the logic gates so that each gate depends only on the output values of the previous gates, then evaluate them one after another. Since topological sorting takes $\mathcal{O}(n)$ time on a graph with $n$ nodes, the CVP takes $\mathcal{O}(n)$ time to solve on a single processor. The CVP embodies the intuition of what an inherently serial problem should be, which is that each step easily depends on the previous step in an easy way, but skipping steps is difficult. As it turns out, any $\mathsf{P}$ problem is efficiently reducible to the CVP, thus making it $\mathsf{P}$-complete.
This gives us an intuitive rule: If the problem appears to have a linear computational graph similar to the CVP, then it is likely inherently serial. Indeed, this intuition guided us in discovering that some problems in Section 4.4 are inherently serial. Furthermore, it is widely believed that $\mathsf{P}$-complete problems are outside of $\mathsf{TC}$, and there is an extensive list of such problems in [34], thus suggesting that inherently serial problems are common and likely to be encountered by an AI.
4.4 Sequential Decision Problems
(a) MCTS in Hex board game: Performance improves with more MTCS expansion nodes across all training regimes. Perfect play is only possible with test-time MCTS. Data from Villalobos & Atkinson [120].
(b) Actor & critic networks’ depth vs. width in locomotion & manipulation. A 2× deeper network (8 layers, width 256) outperforms a 16× wider network (4 layers, width 4,096) in Humanoid, a locomotion task with the largest observation & action spaces among the three. Data from Kevin et al. [53].
Figure 6: Empirical serial scaling on (a) Hex board game and (b) locomotion & manipulation.
As we have seen previously, inherent seriality arises from chains of non-trivial computations—whether rule-based logic or physical interactions—that cannot be compressed into fewer steps. Here, we discuss that the “sequential” nature of sequential decision problems—central to robotics and optimization—also imposes a similar requirements which favor serial solutions over parallel ones.
The goal in sequential decision problems is to obtain an optimal policy $\pi^{*}(s)=\arg\max_{\pi}J(\pi)$, where $J(\pi)$ is the expected return under policy $\pi$ in a Markov Decision Process (MDP) with finite horizon $N$, and discount factor $\gamma\in[0,1]$. Therefore, given a state $s$, the policy outputs an optimal action $a$ from a limited possibilities (discrete action) or with finite precision (continuous action).
Sequential decision problems are inherently serial. In theory, it can be shown that there exists an MDP for which computing the optimal policy is inherently serial. As detailed in Appendix D, the proof relies on a problem that is inherently serial to even approximate, and shows that selecting an optimal action is effectively approximating the problem, which is inherently serial by design.
However, in practice, it remains unclear whether this problem is likely serial. Below, we offer some intuition suggesting that it is. Consider policy gradient methods [113, 125], including popular variants like PPO [98], widely used in applications such as LLM fine-tuning [83]. The gradient is given by
$\nabla_{\theta}J(\pi_{\theta})=\sum_{t=0}^{N-1}\gamma^{t}\;\mathbb{E}_{(s_{t},a_{t})\sim d^{\pi}_{t}}\left[\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\;R_{t}\right],$
where $d^{\pi}_{t}$ is the distribution of $(s_{t},a_{t})$ under policy $\pi_{\theta}$ at time $t$, and $R_{t}=\sum_{k=t}^{N-1}\gamma^{k-t}r_{k}$ denotes the return from time $t$ to the episode’s end.
Given a sufficiently expressive model $\theta$, convergence to the optimal policy via policy gradient requires the return estimate $R_{t}$ to be unbiased [^4] Convergence to suboptimal policies is still possible with biased return estimates [77, 116].. As we will argue, this estimation step represents a key inherently serial bottleneck of sequential decision problems.
Computing the return is likely inherently serial. To compute returns in parallel, not linearly dependent on the trajectory length, it requires accessing state $i$ (and hence obtaining reward $r_{i}$) in fewer than $i$ steps. This is unlikely for real-world MDPs given that CA’s 8-rule transition tables (Section 4.1) or physical interactions (Section 4.2) are sufficient to prevent this. While an exponential-compute parallel algorithm exists [^5] It is possible to random a trajectory and check for its consistence with the current MDP and policy. While this is fast and parallelizable, the trajectory possibilities grow exponentially with the trajectory length $n$., a more efficient way is go through each state in the trajectory one by one serially. Thus, although scaling parallel computation—via aggregating multiple trajectories—can reduce variance in gradient estimation and accelerate convergence [112, p. 93], it cannot replace the accurate return estimation only enabled by serial computation.
This inherent seriality motivates model-based reinforcement learning, where the return is computed by unrolling an internal model step-by-step. For instance, Monte Carlo Tree Search (MCTS)—which increases serial computation via tree expansion and reduces return estimation bias—has achieved superhuman-level performance in diverse board games [103, 104, 97]. Villalobos & Atkinson [120] demonstrate a consistent improvement in the Hex board game from increasing MCTS expansion nodes (see Figure 6(a)). The benefits of increased serial computation are not limited to model-based methods. In model-free settings, Kevin et al. [53] show that deeper (more serial) networks significantly outperform wider (more parallel) ones on locomotion tasks such as Humanoid (see Figure 6(b)).
What does this mean? The roles of serial and parallel computation in reinforcement learning are distinct. Without sufficient serial computation, regardless of the amount of parallel computation, RL will not converge to the optimal policy.
4.5 Math Question Answering
Figure 7: Longer reasoning chains (shown in different colors) vs. majority voting (shown as dots along each line) macro avg. over AMC, MATH, AIME, and Olympiad-bench. Data from Aggarwal & Welleck [2].
Instead of planning through a complex MDP as in sequential decision problems, advanced math question answering exhibits inherent seriality through logical step-by-step reasoning. As shown in Figure 2(E), given a question as input tokens, the model outputs the solution autoregressively—generating each token from a limited set of possibilities. Note that the inherent seriality does not stem from the autoregressive output format; rather, it is a property of the problem that demands serial computation, regardless of the number of output tokens.
Math QA is likely inherently serial. Solving grade-school mathematics, GSM8K [18], which has been used to benchmark reasoning capabilities in LLMs, can be formalized as dependency graphs [131]. A solution is to traverse the graph in topological order and performing necessary arithmetic operations sequentially [131]. This resembles the Arithmetic Circuit Value Problem [34, p. 124], a generalization of the standard CVP to arithmetic operations. Like CVP, this problem is $\mathsf{P}$-complete—that is, inherently serial. Since the inherent seriality pervades even in a relatively simple math QA, it likely extends to more challenging competition-level problems, such as those in AMC, MATH, AIME, and Olympiad-bench.
In mathematics QA, as shown in Figure 7, Aggarwal & Welleck [2] demonstrate that sequential scaling with longer reasoning chains consistently outperforms parallel scaling, which aggregates many shorter chains via majority voting. This pattern holds across mathematical datasets of varying difficulty, including AMC, MATH, AIME, and Olympiad-bench, controlled for the same token budget.
Similarly, science QA also appears to exhibit inherent seriality. Muennighoff et al. [78] report that in GPQA Diamond [90], a PhD-level Google-proof science question set, sequential scaling yields consistent accuracy improvements, only limited by the model’s context window, and is significantly more efficient than parallel scaling (via majority voting), which plateaus.
What does this mean? A certain aspect of complex question answering requires building the answer step-by-step over a computation graph. This aspect, likely shared among math and science QA, requires serial computation.
5 Diffusion model’s computation is not serial
The first part of this paper examined inherently serial problems, which require models capable of serial computation. This section turns to modern ML models and their computational capacity. We provide elaborations on MLPs, Transformers, SSMs, RNNs, and Chain-of-Thought in Appendix B.
Figure 8: The backbone network takes as input a sequence of tokens, and a single noisy output token. This is repeated until the output token is fully denoised.
Here, we focus on diffusion models [42, 108, 107]. These models are widely used in image/video generation  [24, 12], other vision tasks such as depth estimation [96, 27, 52], and language modeling tasks [58, 80, 4].
Consider a problem with input $x_{1},\dots,x_{N}$ and fixed output $x_{N+1}$ (as in Figure 2). A diffusion model uses a backbone neural network $\theta$ which takes the inputs and a noisy version of $x_{N+1}$, denoising it over $T$ steps until it becomes the output (see Figure 8). This models the conditional distribution $p_{\text{truth}}(x_{N+1}\mid x_{1},\ldots,x_{N})$ or concisely $p_{\text{truth}}$.
While the number of denoising steps $T$ is scalable—more denoising steps for finer approximation of $p_{\text{truth}}$, in practice, diffusion models converge rapidly [66]. For instance, image generation plateaus at 300 steps [79], depth estimation shows little difference between 5 and 100 steps [88], and language modeling yields similar perplexity for 32 vs. 1024 steps [5]. With distillation, 1–4 steps suffice without much loss [132, 64, 109, 61, 93]. Such rapid convergence suggests that the effective computation depth of diffusion models is low. It would be surprising if the underlying computation were truly serial.
Diffusion models with a $\mathsf{TC}^{0}$ backbone can only solve problems in $\mathsf{TC}^{0}$. Relying on the results that the backward diffusion converges to $p_{0}$ at the rate $TV=\mathcal{O}(d/T)$ [57, Thm. 1], where $TV$ is the total variation between $p_{0}$ and the denoising $p_{\theta,0}$, and $d$ is the dimension of $x_{N+1}$. For precision $\epsilon>0$ required by a task, there exists a constant $T$ such that, for any input size $N$, the process converges in total variation to within $\epsilon$ at step $T$—effectively solving the task. We arrive at the following theorem:
Theorem 5.1 (Informal).
If a problem can be solved by a diffusion model with a $\mathsf{TC}^{0}$ backbone with high probability with infinite diffusion steps, then the problem itself is in the parallelizable class $\mathsf{TC}^{0}$.
The formal statement and proof of this theorem are provided in Appendix C. It has two parts. The first part uses the fast convergence rate of the Euler–Maruyama method for solving the diffusion equation to show that we do not need many steps to approximately and correctly sample from the diffusion model.
The second part uses a standard derandomization argument to remove the “high probability” qualifier and obtain a deterministic solution. The idea is, for each length $n$, we replicate the network $k(n)$ times. Each network must take one seed. For each choice of $k(n)$ seeds $s_{1},\dots,s_{k(n)}$, we have a particular deterministic model:
$x\mapsto\text{majority}(f(x,s_{1}),f(x,s_{2}),\dots,f(x,s_{k(n)})).$
If $s$ is randomly sampled, then let the probability that $f(x,s)$ is wrong be upper-bounded by a constant $p<1/2$. By Hoeffding’s inequality [43], the probability that the majority vote is correct is $\geq 1-e^{-2k(n)(p-1/2)^{2}}$.
Sample $k(n)$ random seeds, and fix them. This provides a deterministic model. Now, we try this deterministic model on every single possible input of length $n$. There are only #$\#\text{vocab}^{n}$ of them. If we set $k(n)\geq\frac{\ln(\#\text{vocab})}{2(p-1/2)^{2}}$, then by the union bound, the probability that the majority vote is correct on all inputs of length $n$ is nonzero. Thus, there exists a specific choice of random seeds $(s_{1},\dots,s_{k(n)})$ that makes the compound model correct on all inputs of length $n$.
This construction is nonuniform precisely in the part where we have only shown the choice of seeds exists. To actually find these seeds may take exponential time.
What does this mean? Diffusion models only provide a constant amount of additional serial computation. The above theorem precludes the use of diffusion models as a scalable means of increasing serial computation. Unlike Chain-of-Thought (CoT), which genuinely adds serial compute (see Appendix B), diffusion models do not.
Limitation. The theorem doesn’t apply: (1) If the dimension of $x_{n+1}$ is increasing. However, diffusion language modeling, as output length increases, has yet to show promising results [5, 65, 37, 92]. (2) If the backbone is either poorly trained or trained under a different objective. This leaves open the possibility of serial-compute-scalable generative models beyond score-based diffusion.
6 Related Works
Recent work on the formal theory of neural networks highlights a recurring observation: many models that are parallelizable struggle to solve inherently serial problems [41, 111, 9, 39, 137, 94]. This has led to the “parallelism tradeoff” hypothesis—that all parallelizable models, irrespective of design, must necessarily fail to solve inherently serial problems [70, 71].
Our position builds on the classical “depth-width tradeoff” in the theory of parallel algorithms [121], which distinguishes total compute from serial compute via the notions of “work” and “depth” by Blelloch [10]. This distinction underlies Amdahl’s law [3, 38], as inherent seriality imposes lower bounds on achievable speedups.
We also draw from the concepts of $\mathsf{P}$-completeness [34, Ch. 8] and computational irreducibility [126], which characterize systems whose behavior cannot be efficiently reduced. While these ideas are well-established in complexity theory, our work brings them into the machine learning context—extending them to real-world tasks such as sequential decision-making and question answering, and highlighting the gap between these tasks and the limited serial capabilities of modern machine learning models.
7 Limitations
Our conclusions rely on the widely-held but unproven assumption that $\sf{TC}\subsetneq\sf{P}$ (2.3). If this assumption is overturned, the serial/parallel dichotomy would be invalid. Moreover, our theoretical arguments apply only to the general case; although the intuition may extend to practical settings, such generalization is not guaranteed. In particular, the average-case complexity [11] of many problems could differ or even allow parallel solutions. For especially hard problems (e.g., $\mathsf{NP}\setminus\mathsf{P}$ and beyond), exponential cost—rather than serial depth—may dominate as the primary bottleneck. Thus, our focus is on real-world problems of practical difficulty.
Our theorem on diffusion models applies only when the output dimension remains fixed. If it grows with problem size, the result may not hold—though current empirical evidence in language modeling does not suggest strong serial scaling. The theorem also assumes a well-trained backbone using a score-matching objective, and may not apply to poorly trained models, especially early in training.
Finally, beyond theoretical and circumstantial evidence, more empirical work is needed to quantify the degree of seriality in practice—particularly the benefits of increased serial compute. A promising direction is to benchmark real-world tasks under varying ratios of serial and parallel computation.
8 Acknowledgements
We thank William Merrill and Anant Sahai for their insightful feedback and technical corrections, and Yossi Gandelsman and Yuatyong Chaichana for feedback that helped improve the presentation of this paper.
Appendix A $\mathsf{TC}^{0}$ and $\mathsf{TC}$ Classes
Definition A.1.
Let $i\in\mathbb{N}$. A decision problem $L\subseteq\{0,1\}^{*}$ is in the complexity class $\mathsf{TC}^{i}$ if there exists a family $\{C_{n}\}_{n\in\mathbb{N}}$ of Boolean circuits such that:
  • Each circuit $C_{n}$ decides whether $x\in L$ for all $x\in\{0,1\}^{n}$.
  • Each circuit $C_{n}$ has size polynomial in $n$, i.e., $|C_{n}|=\mathcal{O}(n^{k})$ for some constant $k$.
  • Each circuit has depth $\mathcal{O}(\log^{i}n)$.
  • The circuit gates are of unbounded fan-in AND, OR, NOT, and MAJORITY gates. A majority gate outputs 1 if more than half of its inputs are 1.
  • The family $\{C_{n}\}$ is $\mathsf{L}$-uniform, where $\mathsf{L}$ stands for “logspace”. This means that there exists a deterministic Turing machine that, on input $(\underbrace{11\cdots 1}_{n\text{ repeats}},i)$, outputs the $i$th bit of the description of $C_{n}$ using a working tape of length only $\mathcal{O}(\log n)$.
In addition, a decision problem $L\subseteq\{0,1\}^{*}$ is in the complexity class $\mathsf{TC}$ if it is in the union of all classes $\mathsf{TC}^{i}$ over $i\in\mathbb{N}$, that is,
$\mathsf{TC}=\bigcup_{i\in\mathbb{N}}\mathsf{TC}^{i}.$ (1)
In the literature, $\mathsf{TC}$ is also called $\mathsf{NC}$, or “Nick’s Class.”
Define the $\mathsf{L}$-uniform many-one reduction relation $\leq_{m}^{\mathsf{L}}$ as follows: Given two languages $L,L^{\prime}$, that is, two sets of binary strings, we say $L\leq_{m}^{\mathsf{L}}L^{\prime}$ if and only if there exists a function $f$ such that $x\in L$ if and only if $f(x)\in L^{\prime}$, and the function $f$ is computable by a deterministic Turing machine with a logspace working tape.
A decision problem is $\mathsf{P}$-complete if and only if there exists a Turing machine that decides it in polynomial time, and any problem decidable in polynomial time can be $\mathsf{L}$-uniformly many-one reduced to it.
Appendix B Computation capabilities of modern machine learning methods
In this section, we point out an observational fact that machine learning methods that allow parallelization of their computation graph, such as MLPs, Transformers, and SSMs, have so far been shown to lack the ability to solve inherently serial problems. On the contrary, the only methods that are capable of solving serial problems are those that cannot be parallelized, such as RNNs, repeating layers, and Chain-of-Thought (CoT).
In terms of computational complexity, multi-layer perceptrons (MLPs) have been formalized as a family of Boolean circuits with threshold gates called $\sf{TC}^{0}$ [85]. Consider a family of MLPs with constant $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ neurons per layer, and $\mathcal{O}(1)$ numerical precision. A problem is in $\sf{TC}^{0}$ if and only if it is decidable by one forward pass through an MLP under this setup.
It is widely suspected that $\sf{TC}^{1}$ is strictly larger than $\sf{TC}^{0}$. Therefore, any $\sf{TC}^{1}$-hard problem requires computations deepening at a rate of at least $\mathcal{O}(\ln n)$. We may consider the problem as being serial in a weaker, logarithmic sense. A simple example of $\sf{TC}^{1}$ is the word problem of the symmetric group on 5 elements $S_{5}$: Given $g_{1},g_{2},\dots,g_{n}\in S_{5}$, find $\prod_{i}g_{n}$ [63]. Intuitively, this is because $S_{5}$ is not a solvable group, and therefore there is no better way to multiply its elements than via a binary tree, which has $\mathcal{O}(\ln n)$ layers. Attempting to perform this with just $\mathcal{O}(1)$ layers would effectively require one to memorize the entire $S_{5}^{n}\to S_{n}$ multiplication table, which scales exponentially as $e^{n\ln(120)}$, an exponential depth-width trade-off.
To solve the word problem of $S_{5}$ by a Transformer in one forward pass, we present it with input tokens $g_{1},g_{2},\dots,g_{n},y$, where $y$ is a special token. The Transformer’s output at $y$ is read out as the answer. Similarly, we may solve such a problem by presenting the same input tokens to a State Space Model (SSM) or a Recurrent Neural Network (RNN). Since a Transformer with $\mathcal{O}(1)$ layers and $\mathsf{poly}(n)$ dimensions can only solve problems in $\sf{TC}^{0}$ [71] in one forward pass, it cannot solve the word problem of $S_{5}$ that lies outside $\sf{TC}^{0}$.
On the contrary, for an RNN, we can write the multiplication table of $S_{5}^{2}\to S_{5}$ directly into its weights, so it can solve the problem by unrolling for $\mathcal{O}(n)$ recurrence steps with $\mathcal{O}(1)$ layers and $\mathcal{O}(1)$ dimensions. Intuitively, the hidden states of an RNN keep track of the progress of multiplication as it performs the forward passes. However, RNN’s recurrence state dependency renders it non-parallelizable. Several families of SSMs were developed as a compromise that still have recurrence on hidden states, like an RNN, while making forward passes parallelizable, like a Transformer. The prototypical SSM architecture is Mamba [35], though there are many others.
Unfortunately, it has been proven that there is not yet a “free serial-compute lunch,” in the sense that the main families of SSMs proposed so far still cannot solve the word problem of $S_{5}$ under the constraints of $\mathcal{O}(1)$ layers, $\mathsf{poly}(n)$ dimensions per layer, and one forward pass. Despite its apparent recurrence, the hidden state offered by an SSM is weaker in a computational sense than that offered by an RNN [110, 70, 74]. This theoretical fact rhymes with the empirical fact that in practice, MLPs, Transformers, and SSMs are all more parallelizable than RNNs.
Only a genuinely serial method has been shown to go beyond the $\mathsf{TC}^{0}$ class. In addition to RNNs, this includes repeating layers and Chain-of-Thought (CoT). Repeating layers for $\mathcal{O}(\ln n)$ times enables a standard Transformer to solve tasks in the $\mathsf{TC}^{1}$ class [72]. With $\mathsf{poly}(n)$ CoT, requiring multiple forward passes before producing a final answer, a Transformer can be lifted from $\mathsf{TC}^{0}$ to $\mathsf{P}$ [31, 59]. As discussed in Section 4.5, the power of CoT has been well-attested in practice by the improved performance of reasoning models in complex math and science tasks.
Such a uniformity restriction is necessary for technical reasons. Specifically, it is necessary because one may hide a large amount of computation into a small circuit that requires a long time to find. The final circuit produced might run in time $\mathcal{O}(\log n)$, but if the time required to find such a circuit requires time $2^{\mathcal{O}(n)}$, then this would not be parallel—in the sense used throughout this paper.
Appendix C Diffusion is in non-uniform $\mathsf{TC}^{0}$
In this section, we consider the non-uniform $\mathsf{TC}^{0}$ class, in contrast to the usual uniform classes. Note that uniform $\text{uniform }\mathsf{TC}^{i}\subseteq\text{non-uniform }\mathsf{TC}^{i}$. We prove that many diffusion models are restricted within that class. We need to assume non-uniformity because, at a certain point in the proof of the main theorem, we merely prove that something exists, without showing that it is also efficiently computable. We will highlight this in the proof.
C.1 Language modeling
At the most abstract level, a language is simply a set of words made of letters. Formally:
  • An alphabet $\Sigma$ is a finite nonempty set. Each element in the alphabet may be called a letter or a token.
  • A word in an alphabet $\Sigma$ is a finite sequence of elements of $\Sigma$.
  • A language $L$ in an alphabet $\Sigma$ is a set of words in the alphabet $\Sigma$.
A prefix language modeling problem is, given a sequence of tokens $x_{1},\dots,x_{n}$, to compute the next token $x_{n+1}$. This is a deterministic formalization of “next-token prediction,” the dominant paradigm in language modeling since the GPT-2 of 2019. This is also a decision problem formulation: input of size $n$, $x_{1},\dots,x_{n}$, and a decision of either “yes” or “no,” $x_{n+1}$.
While usually, a diffusion model is used for generating from a continuous state space such as $\mathbb{R}^{d}$, it can be used to model discrete distributions as well. This is necessary for language modeling. We consider the case closest to continuous state space modeling—quantization: One divides the continuous state space $\mathbb{R}^{d}$ into regions and assigns a token to each region. This then allows sampling a discrete distribution from a diffusion model with continuous state space. Formally, if $\Sigma=\{a_{1},a_{2},\dots,a_{M}\}$ is the alphabet, then we divide $\mathbb{R}^{d}$ into $M$ regions $V_{1},\dots,V_{M}$, such that each region $V_{i}$ maps to a token $a_{i}$.
In non-uniform circuit complexity, for each task instance $n$, we have a $\mathsf{TC}^{0}$ score-network $f_{\theta,n}$ such that:
  • Each $f_{\theta,n}$ takes as input $n+2$ elements $x_{1},\dots,x_{n},x,t$, and produces an output $f_{\theta,n}(x,t|x_{1},\dots,x_{n})$.
  • The family $f_{\theta,n}$ has $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ width, and $\mathcal{O}(1)$ precision.
Comment. The theorem holds for any family of score-networks for which a single forward pass is in $\mathsf{TC}^{0}$. This includes, for example, Transformers and state-space models [71, 74].
Finally, since a diffusion model may solve a problem only with high enough probability, instead of solving it deterministically, we make the following definition: A prefix language modeling problem is solved with constant probability bound if there exists some $\epsilon>0$, such that for each input token sequence $x_{1},\dots,x_{n}$, let $x_{correct}$ be the correct response, then
$p(x_{correct}|x_{1},\dots,x_{n})>p(x^{\prime}|x_{1},\dots,x_{n})+\epsilon,\quad\forall x^{\prime}\neq x_{correct}.$ (2)
C.2 Main theorem
Theorem C.1.
Suppose there exists a $\mathsf{TC}^{0}$ family of score-networks $f_{\theta,0},f_{\theta,1},\dots$, such that for each $n$ and each $x_{1},\dots,x_{n}$, the function $f_{\theta,n}(x,t|x_{1},\dots,x_{n})$ exactly computes the score function of some distribution $\rho_{x_{0}|x_{1},\dots,x_{n}}$ with bounded first moment: $\mathbb{E}_{x_{0}\sim\rho_{x_{0}|x_{1},\dots,x_{n}}}[\|x_{0}\|]\leq 1$.
If this family solves a prefix language modeling problem at the limit of infinite time Score-Matching with Langevin Dynamics (SMLD) [107] with constant probability bound, then the problem is in the $\mathsf{TC}^{0}$ class.
Proof.
The idea of the proof is simple. We first construct a universal $\mathcal{O}(1)$-bound on how many steps are sufficient for sampling the SMLD within a constant probability bound, then we derandomize it according to [40], while still remaining within the $\mathsf{TC}^{0}$ class.
Since the score-network exactly computes the score function for $\rho_{0}$, the score-matching error is exactly zero.
By [57], there exists a constant $c>0$ that does not depend on any parameter in the statement of the theorem (a "universal constant"), such that there exists a noise schedule for DDPM that takes $T$ steps of the following kind. Using the schedule and the score-matching function $f_{\theta}$, we sample from a distribution $\rho_{DDPM,T}$. It satisfies the inequality
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq c\frac{d(\log T)^{3}}{T}$ (3)
Now, since $c$ is a universal constant, we can take $T$ to be $cc^{\prime}d^{2}$, to obtain an upper bound
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq\frac{(\log(cc^{\prime}d^{2}))^{3}}{c^{\prime}d}$ (4)
The key is that $T$ does not increase with $n$.
Since the growth of $\log^{3}$ is dominated by linear growth, for any $\epsilon^{\prime}>0$, there exists a large enough universal constant $c^{\prime}$, such that
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq\epsilon^{\prime}$ (5)
for all $d=1,2,3,\dots$.
Let $\epsilon$ be the constant probability bound, then setting $\epsilon^{\prime}=\epsilon/2$, we find that $\rho_{DDPM,T}$ already solves the problem with constant probability bound.
Now we can derandomize this family, obtaining a $\mathsf{TC}^{0}$ family of Boolean circuits that solves the problem deterministically. The details of the derandomization method appear in [40, Prp. 4.2]. The big picture is as follows: we remove the probability by hard-coding a magic constant [^6] This is the part where we assume non-uniformity. In complexity theory, this is an “advice string”. per member of the family, such that sampling for polynomially many times, and taking the majority output, would always give the correct output. By Hoeffding’s inequality, such magic constants exist for a large enough polynomial.
Comment. The requirement for exact score-matching is necessary for the following two reasons:
First, the full form of the inequality from [57] is
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq c\frac{d(\log T)^{3}}{T}+c\epsilon_{\text{score}}\sqrt{\log T},$ (6)
where the term $\epsilon_{\text{score}}$ denotes the score-matching error between the true score function of $\rho_{\hat{x},0}$ and the approximation $f_{\theta}$. As this extra term increases with $T$, the proof above does not apply.
Second, if we have no requirement on score-matching, then there is essentially no constraint on the computational power of SMLD, since any Turing machine can be constructed in SMLD with some (extremely non-score-matching) network. [^7] An explicit construction is published on Yuxi’s personal website. Yuxi would have published it on arXiv, but arXiv has put the paper “on hold” for months with no sign of actually publishing it.
Practically relevant score-networks are intermediate between two extreme cases. We believe that if $f_{\theta}$ is a good enough, but not perfect, score-matching network, then a generalized version of the above theorem still applies. However, finding the right way to quantify the goodness, as well as proving such a generalization, is left as future work.
Appendix D Inherently Serial Problems in RL
Throughout this section, by a parallel algorithm, we mean specifically an $\mathsf{L}$-uniform $\mathsf{TC}$ Boolean circuit family—as usual throughout this paper.
In this section, we begin with a problem from computational complexity theory that is proven to be impossible to parallelize (assuming, as always, that $\mathsf{TC}\neq\mathsf{P}$), then convert it into a deterministic decision problem. We then prove a theorem, showing that any parallel decision rule for this problem has arbitrarily bad worst-case performance. As special cases, this includes maximizing parallel value functions, maximizing parallel Q-functions, parallel policies, and parallel learning rules that produce parallel policies.
D.1 Definitions
A Boolean circuit $C$ is alternating when it consists solely of $\operatorname{AND}$ and $\operatorname{OR}$ gates such that every $\operatorname{AND}$ gate connects to only $\operatorname{OR}$ gates and every $\operatorname{OR}$ gate connects to only $\operatorname{AND}$ gates.
Alternating circuits are monotonic in the following sense. Consider an alternating circuit that takes $n$ inputs, and let $x,x^{\prime}\in\{0,1\}^{n}$ be two possible inputs to it. If $x\leq x^{\prime}$ (i.e., $x^{\prime}$ is obtained by flipping some zeros of $x$ to ones), then $C(x)\leq C(x^{\prime})$. More generally, for any gate output $C_{i}$ of $C$, we have $C_{i}(x)\leq C_{i}(x^{\prime})$; intuitively, this can be visualized as “hot” wires carrying True-signals monotonically “upwards” through the circuit.
The depth of a gate is the length of the longest directed path from any circuit input to that gate. A chain of the form $\operatorname{OR}\!\rightarrow\!\operatorname{AND}\!\rightarrow\!\operatorname{OR}\!\rightarrow\!\cdots$ therefore has gate depths $1,2,3,\dots$ in order.
For an alternating circuit $C$ on input $x$, the depth-of-1 (DO1) of this circuit configuration is
$d_{1}(C,x)\;:=\;\max\bigl{\{}\text{depth}(g)\;:\;g\text{ outputs }1\text{ on }x\bigr{\}}.$
The problem of computing the DO1 of any circuit configuration is the DO1 problem. Assuming that $\mathsf{TC}\neq\mathsf{P}$, as usual in the paper, then the DO1 problem cannot be solved by a parallel algorithm. In fact, much more can be said. Not only would it be non-parallelizable, any approximation of it is also non-parallelizable.
Fix constants $0<\epsilon<b$. Given $(C,x)$, the problem of computing a number $d(C,x)$ satisfying
$d(C,x)\;\in\;[\epsilon\,d_{1}(C,x),\;b\,d_{1}(C,x)].$
Because a solution for $[\epsilon,b]$ yields one for $[\epsilon/b,1]$, it suffices to consider the special case $b=1$, which we call the $\epsilon$-approximate DO1 problem.
D.2 Non-parallelizability Results
We quote the following result from Kirousis and Spirakis [54].
Theorem D.1.
For every $\epsilon\in(0,1)$, the $\epsilon$-approximate DO1 problem is $\mathsf{P}$-complete. Consequently, if $\mathsf{TC}\neq\mathsf{P}$, no parallel algorithm can solve the $\epsilon$-approximate DO1.
To import the theorem from computational complexity theory into RL theory, we need to construct a specific decision environment in which an agent must perform actions to maximize rewards.
The idea of the following construction is that an approximately optimal agent can be exploited as a problem-solving resource for other ends. Specifically, we will construct some environments in which the first action of the agent is a forced choice between two circuits. In the language of psychologists, we construct two-alternative forced choice experiments. The agent’s choice can then be interpreted as an agent’s judgment as to which circuit has a deeper depth-of-1.
Given a circuit configuration $(C,x)$, we construct, in parallel, many other circuit configurations $(C^{\prime}_{1},1),(C^{\prime}_{2},1),(C_{3}^{\prime},1),\dots,(C_{|C|}^{\prime},1)$, such that $d_{1}(C^{\prime}_{1},1)=1,d_{1}(C^{\prime}_{2},1)=2,d_{1}(C_{3}^{\prime},1)=3,\dots,d_{1}(C_{|C|}^{\prime},1)=|C|$. Then, we perform in parallel all forced choices for these pairs: $(C,x)$ vs $(C^{\prime}_{1},1)$, $(C,x)$ vs $(C^{\prime}_{2},1)$, …, $(C,x)$ vs $(C^{\prime}_{|C|},1)$. At some point, the agent’s forced binary choice should switch from preferring $(C^{\prime}_{k},1)$ to preferring $(C,x)$. This can be taken as a judgment that $d_{1}(C,x)\approx k$. If the agent is approximately optimal, then $d_{1}(C,x)\approx k$ is correct, in the sense that we can guarantee $k\in[\epsilon d_{1}(C,x),bd_{1}(C,x)]$ for some constants $0<\epsilon<b$ that do not depend on either $C$ or $x$.
This is the essential idea of the construction and the subsequent proof. The rest are tedious details designed to close loopholes.
We define the DO1 environment as follows. Given a circuit configuration $(C,x)$, define a corresponding deterministic decision problem as follows.
  1. 1.
    At $t=0$, the agent observes two circuit configurations: the original $(C,x)$ and a length-$k$ alternating chain of the form $\operatorname{OR}\!\rightarrow\!\operatorname{AND}\!\rightarrow\!\operatorname{OR}\!\rightarrow\!\cdots$, whose input is 1.
  2. 2.
    The agent selects one gate from one of the circuits; the unchosen one is then removed. This is the two-alternative forced choice.
  3. 3.
    Subsequently, the agent may select at most one additional gate per time-step, for $H:=\max\{k,|C|\}$ steps, designed so that the agent has enough time to choose every desired gate.
  4. 4.
    The episode ends at $t=H$. If every chosen gate outputs 1, then the reward at this step is $r_{H}=$ the maximal depth among all the chosen gates. Otherwise, $r_{H}=0$. The reward at all other steps is zero.
This means that each DO1 environment can have the following states:
  • No gates are selected.
  • Some gates of $C$ are selected.
  • Some gates of the alternating chain are selected.
Theorem D.2 (optimal decision in the DO1 environment is inherently serial).
Assume $\mathsf{TC}\neq\mathsf{P}$.
  1. (a)
[…]
  2. (b)
    For any parallel algorithm producing a value function $V$ and any $\epsilon\in(0,1)$, there exists a DO1 environment where the greedy policy $\pi_{V}(s):=\arg\max_{a}V(s^{\prime})$ achieves terminal reward $r_{H}<\epsilon r^{*}$, where $r^{*}$ is optimal. Here, $s^{\prime}$ denotes the next state if, at state $s$, the agent performs action $a$.
  3. (c)
    Statement (b) extends to any parallel algorithm that, given a state in a DO1 environment, outputs an action.
Proof.
  1. (a)
    We argue by contradiction. Assume there exists a parallel algorithm that, for some constant $10<\epsilon<$ 1, outputs a value function $V$ satisfying
[…]
    Since any factor $b>1$ can be removed by division, we set $b=1$ without loss of generality.
    Special case. We need to check first the special case where $d_{1}(C,x)=0$. This is done as follows. First, perform a topological sort of $C$, which is parallelizable (in $\mathsf{TC}^{2}$, in fact [21]). This then allows us to find all the gates that are in the first layer. Next, for each gate in the first layer, we test in parallel whether that gate outputs 1. Let their outputs be $y_{1},\dots,y_{m}$. If $\mathrm{OR}(y_{1},\dots,y_{m})=0$, then $d_{1}(C,x)=0$, because alternating circuits are monotonic. Otherwise, $d_{1}(C,x)\geq 1$.
    Having thus handled the special case, we assume that $d_{1}(C,x)\geq 1$ for the rest of the proof.
    Two-alternative forced choice experiments in parallel. Given an input $(C,x)$ with $n:=|C|$ gates and depth-of-1 equal to $d_{1}=d_{1}(C,x)$, choose $m\in\mathbb{N}$ such that $2^{m-1}<n\leq 2^{m}$. For each index $\ell\in\{0,\dots,m\}$ construct, in parallel, a DO1 environment $E_{\ell}$ whose second circuit is an alternating chain of length $2^{\ell}$, with the only input being  1. Within each $E_{\ell}$ evaluate (again in parallel)
    • $V\!\bigl{(}s_{g}^{(\ell)}\bigr{)}$, where $s_{g}^{(\ell)}$ is the state after initially selecting gate $g\in C$;
    • $V\!\bigl{(}s_{\text{chain}}^{(\ell)}\bigr{)}$, where $s_{\text{chain}}^{(\ell)}$ is the state after selecting the input gate of the chain.
    The whole procedure requires
    $(n+1)(m+1)=\bigl{(}|C|+1\bigr{)}\bigl{\lceil}\log_{2}|C|\bigr{\rceil}$
    parallel evaluations.
    Identifying the switchover index. For $\ell=m$ the chain depth $2^{m}$ is at least as long as what depth-of-1 that $C$ can create, hence, if the value function is any good, it should satisfy $V\!\bigl{(}s_{\text{chain}}^{(m)}\bigr{)}\geq V\!\bigl{(}s_{g}^{(m)}\bigr{)}$ for all $g$. As $\ell$ decreases the chain shortens; eventually picking $C$ should be better than picking the chain. This intuition allows us to define the following decision procedure.
    Let $k$ be the smallest $\ell$ such that
    $V\!\bigl{(}s_{\text{chain}}^{(\ell)}\bigr{)}\;\leq\;\max_{g}V\!\bigl{(}s_{g}^{(\ell)}\bigr{)}.$
    Output $d^{\prime}:=2^{m-k}$ as the estimate for $d_{1}$.
    There are two special cases to handle. If the agent rejects the chain even for $\ell=0$, then output $|C|$. If the agent always picks the chain, then output $d^{\prime}=1$.
    Quality of the estimate. Suppose the switchover occurs, such that the agent picks the chain for $E_{k}$, but switches to picking the circuit for $E_{k-1}$. Consider in detail what happens for $E_{k}$.
    Choose $g^{*}$ attaining the maximum in the definition of $k$ and set
    $s:=s_{g^{*}}^{(k)},\quad s^{\prime}:=s_{\text{chain }}^{(k)}$
    by construction, $V(s)\geq V(s^{\prime})$, and by the assumed guarantee on $V$,
    $V(s)\in[\epsilon d_{1},\,d_{1}],\qquad V(s^{\prime})\in[\epsilon d^{\prime},\,d^{\prime}].$
    Hence $d_{1}\geq\epsilon d^{\prime}$.
    Similarly, the argument in the case for $E_{k-1}$ shows $\epsilon d_{1}\leq 2d^{\prime}$, so altogether
    $d_{1}\;\in\;\bigl{[}\epsilon d^{\prime},\,\tfrac{2}{\epsilon}\,d^{\prime}\bigr{]}.$
    A similar argument applies for the two special cases where no switchover occurs.
    Therefore the procedure is an approximation algorithm for DO1 that operates in parallel, contradicting Theorem D.1.
  2. (b)
    A special case of (c).
  3. (c)
    Assume for contradiction that there exists a parallel decision algorithm that is within an $\epsilon$ of optimality. That is, when the decision algorithm is applied in any DO1 environment, it always achieves reward at least $\epsilon r^{*}$. Then, for any particular DO1 environment, its first action must be an approximate solution as to whether $(C,x)$ or the chain possesses larger depth-of-1. Now, the same construction of (a) yields a parallel algorithm solving $\epsilon^{\prime}$-approximate DO1, again contradicting Theorem D.1.
Comment. By prepending beneath the circuit $C$ a length-$|C|$ alternating chain of gates, and making some minor adjustments to the proof, we can show that any parallel decision algorithm achieves linear regret in the worst case, meaning that $r_{H}-r^{*}=\Theta(H)$.
Although our analysis uses the DO1 problem in particular, there are some other $\mathsf{P}$-complete problems, such as linear programming, that are $\mathsf{P}$-complete to approximate. See for instance [91, 99, 25, 34] for some examples. This leads us to conjecture that inherent seriality may be a fairly common phenomenon in RL, not particular to the DO1 setup.
The barrier in Theorem D.2 can be bypassed in several ways:
  • In the unlikely case that $\mathsf{TC}=\mathsf{P}$ is proven, it would have great consequences for complexity theory in general, analogous to the case where $\mathsf{P}=\mathsf{NP}$ is proven. This rejects the fundamental assumption in the theorem.
  • By employing a serial learning algorithm that runs in polynomial (not polylog) time, one may discover the right policy, which can then be compressed down into a policy that runs faster than serial. This bypasses the “$\mathsf{L}$-uniform” part of the obstacle.
  • By allowing the learned policy or value function to be non-parallel, the “$\mathsf{TC}$ circuit family” part of the obstacle is bypassed. This is true for certain RL algorithms, especially model-based methods that simulate the environment dynamics before making a decision.
  • For some RL applications, one may be able to tolerate arbitrarily poor worst-case performance, as long as they occur rarely.
Part (b) of the theorem may explain the observation in Kevin et al. [53]. In that work, they trained both policy networks and value function networks by actor-critic methods, and noted that using deeper networks on both policy and value function improved performance. Part (b) of the theorem suggests that, when the network for approximating the value function contains less serial compute than the environment demands, then the corresponding exact $V$-maximizing policy would suffer arbitrarily bad worst-case performance. Though the theorem does not exactly apply when the policy is inexact, it suggests that the same phenomenon happens for an actual policy network trained to merely approximate the $V$-maximizing policy.